# A General-Purpose Multimodal Foundation Model for Dermatology

### v1

@@ Results
We evaluated PanDerm’s potential to identify digital biomarkers for melanoma progression using dermoscopic images, a relatively underexplored but clinically valuable area [13, 14, 49] (Fig 3f). The evaluation used the CombinMel dataset, an international multi-center cohort with 680 dermoscopic images of invasive primary melanoma from 370 patients. PanDerm’s performance was assessed for predicting melanoma metastasis (Fig 3g, details in Methods). In binary classification (control vs. metastasis), PanDerm achieved an AUROC of 0.964 (95% CI 0.937-0.991), outperforming the next-best model by 2.0% (P = 0.073) (Fig 3e). For the more challenging three-class classification (control vs. local metastasis vs. distant metastasis), PanDerm significantly outperformed the next-best model by 2.8% (P $<$ 0.05) in weighted F1 score (Extended Data Table 6).
To evaluate the prognostic value of PanDerm’s metastasis prediction scores, we conducted survival analyses using Kaplan-Meier analysis and Cox proportional hazards regression. Kaplan-Meier analysis (Fig 3h) demonstrated that patients stratified into the high-risk group based on PanDerm’s prediction scores had significantly shorter recurrence-free intervals (RFI) compared to those in the low-risk group (HR: 5.63, 95% CI: 2.87-11.02, P $<$ 0.001). We then performed multivariate Cox regression, incorporating PanDerm’s prediction scores alongside clinical variables such as sex, age, Breslow thickness, ulceration, dermal mitosis, location, and melanoma subtype. The resulting hazard ratios (Fig 3i) indicated that PanDerm’s metastasis prediction was the strongest predictor of RFI among all variables considered. To further assess predictive accuracy, we evaluated the time-dependent AUC at different intervals, comparing PanDerm’s score with single and multiple clinical variables (Fig 3j). PanDerm achieved AUCs of 0.950 (95% CI 0.910-0.991), 0.931 (95% CI 0.887-0.976), and 0.909 (95% CI 0.880-0.937) at 3, 5, and 7 years, outperforming multi-clinical variables by 6.8%, 2.9%, and 5.0%, respectively. Combining PanDerm’s score with clinical variables further improved AUCs at 5 and 7 years (Fig 3j). PanDerm also outperformed all pretrained models (Fig 3k), exceeding the next-best model (DINOv2) by 2.3%, 3.0%, and 2.5% at 3, 5, and 7 years, respectively. Details on the experimental setup and datasets are provided in Methods.
@@
We evaluated PanDerm’s performance in skin phenotype, risk assessment, and malignant lesion screening using 3D TBP [2, 3, 50] (Fig 4a). Unlike dermoscopy, TBP allows for the integration of broader patient-level data, facilitating the assessment of risk factors such as photodamage and nevus count, both critical for melanoma risk assessment [4, 5, 51]. Using in-house TBP data, PanDerm achieved a weighted F1 score of 0.896 (95% CI 0.879-0.913) for photodamage assessment and an AUROC of 0.983 (95% CI 0.979-0.987) for nevus counting, outperforming all other models (P $<$ 0.05 and P $<$ 0.001, respectively; Fig 4b, c, g). Notably, PanDerm maintained its performance advantage even with limited labeled data, outperforming the next-best model while using only 10% of the training data (Extended Data Fig 4). In lesion-specific risk stratification, PanDerm also ranked first with an AUROC of 0.705 (95% CI 0.698-0.712) and BACC of 0.657 (95% CI 0.6513-0.663), with all results statistically significant (P $<$ 0.001; Fig 4d, h).
@@
Figure 4: Skin phenotype assessment and melanoma screening using TBP. a. Illustration of PanDerm processing multimodal TBP data (HOP & MYM dataset) for total body skin examination, enhancing skin phenotype score quantification, risk stratification, and malignant lesion screening for advanced early melanoma detection. b-c. Class distribution of skin phenotype for solar damage risk ((b)) and nevus count ((c)) in the HOP and MYM datasets. d-e. Class distribution of risk groups and malignant lesions. g. Solar damage risk assessment and nevus counting performance, measured by W_F1 (weighted F1) and AUROC, respectively. h. Risk prediction performance, measured by AUROC and BACC (Balanced Accuracy). Error bars in g-h show 95% CIs; bar centers represent mean value. The estimates are computed with nonparametric bootstrapping using 1000 bootstrap replicates. P-values are calculated with a two-sided t-test. j. Malignant lesion screening performance, measured by sensitivity for malignant lesions. Left: using only Total Body Photography (TBP) data. Right: integrating measurement information. Numbers below each bar indicate the recommended count of suspicious lesions. k. Number of malignant lesions detected in the test set. f. UMAP plot of PanDerm screening results for all test lesions. i. UMAP plot of human screening results for all test lesions. l. UMAP plot of PanDerm risk prediction results for all test lesions.
@@
For skin cancer screening using TBP, PanDerm effectively identified malignant lesions among a large number of benign lesions in a highly imbalanced dataset (216 malignant vs. 197,716 benign lesions) from the HOP [52] and MYM [53] cohort (Fig 4e). PanDerm outperformed the next-best model in sensitivity by 4.2% when using only TBP data (Fig 4j left), and by 3.5% when integrating additional measurement data (Fig 4j right), achieving a sensitivity of 0.893. It successfully detected malignant lesions in 79 out of 80 patients and recommended significantly fewer benign lesions for dermoscopy compared to melanographers (3498 vs. 8913), thereby reducing unnecessary examinations by approximately 60.8% (Fig 4j, k; Extended Data Table 10).
@@
The distribution of lesions, visualized through UMAP plots (Fig 4f), demonstrated that PanDerm’s feature space effectively differentiates suspicious lesions, aligning with the clinical “ugly duckling” concept [54]. This concept involves identifying atypical lesions by comparison with other lesions on the same individual. Clustering patterns (i.e., clusters of lesions with similar characteristics) in PanDerm’s risk stratification (Fig 4l) corresponded closely with those observed in human screenings (Fig 4i), illustrating PanDerm’s exceptional performance in malignant lesion screening. Additional details on the experimental setup, dataset, measurement information and variable importance for TBP, and performance metrics are available in the Methods section, Extended Data Table 7-10, and Extended Data Fig 5.
@@
We evaluated PanDerm’s performance on skin lesion segmentation using two benchmark datasets: ISIC2018 (task 1) [55] and HAM10000 [35]. PanDerm was compared against SL-Imagenet, autoSMIM [34], and BATFormer [34]. PanDerm significantly outperformed all models, surpassing the next-best by 3.1% and 1.9% in Jaccard index (JAC) on ISIC2018 and HAM10000, respectively (P $<$ 0.001; Extended Data Fig 1a, b). Its Dice score (DSC) was also significantly higher than that of all other models across both datasets (P $<$ 0.001). PanDerm’s performance was particularly noteworthy in label-limited scenarios, matching the next-best model while using only 5% of the training data (104 and 350 images for ISIC2018 and HAM10000, respectively; Extended Data Fig 1c, d). To further validate its capabilities, PanDerm was compared with MedSAM [56], a state-of-the-art medical segmentation foundation model, and outperformed it in JAC by 0.5% (P = 0.025 and 0.112; Extended Data Table 13) on the two datasets. This is notable since PanDerm does not use segmentation-specific datasets during pretraining, unlike MedSAM. Additionally, PanDerm uses a smaller input size (224 $\times$ 224) compared to MedSAM (1024 $\times$ 1024), resulting in significantly lower computational costs. On an RTX3090 GPU, PanDerm’s fine-tuning and inference speeds were approximately 5 and 4 times faster than MedSAM’s, respectively (Extended Data Table 14). Further details on segmentation tasks are provided in Methods, with qualitative results in Extended Data Fig 2 and quantitative results in Table 11-14.
Reader study 2: Human-AI collaboration
@@
To assess PanDerm’s clinical applicability, we conducted a reader study to investigate its potential to improve clinicians’ diagnostic accuracy. The study included 41 human raters with diverse experience levels. We compared their accuracy in diagnosing seven classes of pigmented lesion classes using dermoscopic images, both with and without PanDerm’s multi-probability prediction support (Fig 5a). PanDerm’s support significantly increased overall diagnostic accuracy from 0.69 (95% CI 0.65-0.73) to 0.80 (95% CI 0.76-0.84, P $<$ 0.001; Fig 5b). A subgroup analysis revealed that raters with less experience benefited the most, showing accuracy improvements of 17% (P = 0.0082) for those with low experience and 12% (P $<$ 0.001) for those with medium experience, while highly experienced raters showed a 6% improvement (P = 0.039; Fig 5c, Extended Data Table 25). The class-specific analysis demonstrated significant accuracy improvements in 4 out of the 7 classes (P $<$ 0.05; Fig 5d, Extended Data Table 24). For melanoma specifically, PanDerm improved the accuracy of human raters from 0.69 (95% CI 0.64-0.74) to 0.83 (95% CI 0.79-0.87, P $<$ 0.001). Additional details on the reader study can be found in the Methods section.
Figure 5: Performance of PanDerm in a Human-AI collaboration setting. a. Overview of the reader study: 41 users participated, answering a total of 3,200 questions. The maximum, median, and minimum number of questions answered per user were 1,040, 40, and 20, respectively. b. Comparison of diagnostic accuracy without support and with PanDerm support (P $<$ 0.001; two-sided paired t-test; n = 41 readers). c. Comparison of diagnostic accuracy without support and with PanDerm support, grouped by experience level: Low (n = 11 readers), Medium (n = 21 readers), and High (n = 9 readers). d. Comparison of diagnostic accuracy without support and with PanDerm support, grouped by diagnostic classes. Abbreviations: MEL, melanoma; BCC, basal cell carcinoma; AKIEC, actinic keratosis/intraepidermal carcinoma; BKL, benign keratinocytic lesion; NV, melanocytic nevus; DF, dermatofibroma; VASC, vascular lesion.
@@ Discussion
Despite significant advances in AI technology, its application in clinical medicine remains fragmented and underutilized. Current AI systems are often restricted to isolated tasks, unable to address the diverse demands of medical decision-making. This limits AI’s potential in supporting clinicians in disease diagnosis and management. Dermatology, with its complex requirements including holistic patient assessment, lesion-specific analysis, and potential use of various imaging modalities, serves as an ideal use case for demonstrating AI’s capabilities across multiple interconnected clinical tasks. Success in this domain could pave the way for broader adoption of AI models across healthcare.
@@
To assess PanDerm’s potential for extensive clinical applications, we evaluated the foundation model across 28 benchmark datasets encompassing the most common dermatology-related clinical tasks. PanDerm outperformed existing models in skin cancer-related tasks, including risk stratification of individual phenotypes and lesions, detection of lesion changes and malignancy, and metastasis prediction and prognosis. Moreover, it demonstrated superior performance in diagnosing neoplastic and inflammatory skin diseases. Notably, PanDerm achieved these results often using significantly less training data (5-10%) than existing models, addressing the critical challenge of acquiring high-quality labeled data that typically require specialist input. Reader studies further validated PanDerm’s real-world applicability, providing evidence of improved diagnostic accuracy with AI support. Collectively, these findings underscore PanDerm’s transformative potential, enhancing AI integration into clinical workflows.
The scaling behavior observed in PanDerm’s performance, particularly in relation to pretraining dataset size, aligns with recent trends in foundation model development [20, 22, 23]. However, our study uniquely demonstrates this phenomenon in dermatological AI, where large-scale data acquisition poses distinct challenges. The positive impact of increasing dataset size and diversity on model performance underscores the importance of collaborative data collection efforts across institutions and countries. Our comparative analysis also provides key insights into the effectiveness of different self-supervised learning approaches. Using CLIP [37] as a teacher model for semantic feature alignment resulted in superior training data efficiency (Fig 1f), significantly outperforming DINOv2 [40]. This is especially relevant in healthcare, where assembling datasets comparable to those used in natural image processing is often impractical. For context, DINOv2 typically requires massive datasets like LVD-142M (142 million images), a scale that is rarely achievable in medical domains. Furthermore, our masked feature reconstruction approach demonstrated a stronger capacity for capturing semantic-level information compared to methods like MAE [19], which is a critical capability in the diagnosis of skin diseases, where subtle semantic features can be diagnostically significant. These advantages allowed PanDerm to consistently outperform traditional models such as ResNet50 [58] and ViT large [38], as well as recent generalist medical foundation models such as Med-PaLM-M [59] and BiomedGPT [60] across a broad spectrum of clinical tasks (Extended Data Table 30). The superior performance of PanDerm over generalist medical AI models can be attributed to several factors. Despite their broad scope, current generalist models struggle to achieve clinical-grade performance in specialized medical domains. This underperformance stems from limitations such as insufficient domain-specific data, lack of coverage for specialized modalities like TBP, and challenges in acquiring large, high-quality datasets across diverse medical domains. Additionally, potential modality conflicts [61] and negative transfer issues [60] between distinct medical domain data further hinder their performance in highly specialized clinical tasks. In contrast, PanDerm’s focused approach to skin diseases overcomes these limitations, resulting in more clinically relevant performance.
Furthermore, PanDerm demonstrated robust performance across diverse demographic factors, including different body locations, age groups, genders, and skin tones, setting it apart from previous models (Extended Data Table 21-22). Notably, minimal bias was observed across different skin tones, contrasting sharply with prior studies [62, 63, 64] that reported significant disparities in AI performance for dermatological conditions among diverse populations. This equitable performance, attributed to PanDerm’s diverse training data and semantic-focused approach, has important implications for reducing health disparities and promoting equitable healthcare delivery in dermatology. PanDerm’s demonstrated clinical applicability in real-world scenarios marks a significant step towards integrating AI support into clinical practice. The model showed considerable promise in enhancing early melanoma detection, facilitating melanoma screening, and improving diagnostic accuracy when used in collaboration with clinicians. These capabilities align closely with the vision of AI as an augmentative tool in healthcare [65, 66, 67].
Despite these promising results, our study has limitations that warrant further investigation. Currently, PanDerm focuses primarily on imaging data and tasks related to skin cancer diagnosis. Integrating large language models [68, 29] and additional data modalities, such as medical reports, proteomics, genomics, and microarray data, into PanDerm’s training could further enhance its diagnostic capabilities and comprehensive understanding of complex medical information. Additionally, while our training data is diverse, its geographic concentration may not fully represent global variation. Future work should prioritize expanding the geographic and ethnic diversity of the training dataset to improve global applicability. In the future, developing foundation models for skin disease diagnosis should emphasize expanding data collection through international institutional collaborations, such as those led by the International Skin Imaging Collaboration (ISIC) [57], while ensuring careful consideration of data quality and potential biases in large-scale sources. Such efforts will be crucial for advancing dermatological AI and enhancing its clinical utility across diverse populations worldwide.
In conclusion, PanDerm represents a significant advancement providing a robust and versatile multimodal foundation model for the diagnosis of skin diseases. Its development addresses key challenges in the field, including data curation, model architecture, training strategy, and clinical validation. By leveraging a large-scale, diverse dataset and advanced self-supervised learning techniques, PanDerm demonstrates superior performance across a broad range of dermatological tasks, consistently outperforming all comparative pretrained models. The principles and methodologies established in this study could serve as a roadmap for developing foundation models in other medical specialties, potentially accelerating the integration of AI support in healthcare.

### v2

@@ Results
We explored PanDerm’s potential to predict melanoma progression from dermoscopic images, an emerging approach that could provide valuable prognostic information at the time of diagnosis [13, 14, 46] (Fig 3e). We evaluated this capability using 680 dermoscopic images from 370 patients with invasive primary melanoma across multiple international centers (Fig 3f). PanDerm demonstrated exceptional accuracy in distinguishing melanomas likely to metastasize, achieving an AUROC of 0.964 (95% CI 0.937-0.991), surpassing the next-best model by 2.0% (P = 0.073) (Fig 3d). It also showed strong capability in differentiating between local and distant metastasis, outperforming existing methods by 2.8% (P $<$ 0.05) in weighted F1 score (Extended Data Table 13).
To validate PanDerm’s clinical utility for patient risk stratification, we conducted survival analyses using Kaplan-Meier analysis and Cox proportional hazards regression. Patients classified as high-risk by PanDerm showed significantly shorter recurrence-free intervals (RFI) compared to those in the low-risk group (HR: 5.63, 95% CI: 2.87-11.02, P $<$ 0.001) (Fig 3g). When compared alongside standard clinical risk factors - including sex, age, Breslow thickness, ulceration, dermal mitosis, location, and melanoma subtype - PanDerm’s predictions emerged as the strongest indicator of recurrence risk in multivariate Cox regression (Fig 3h). It maintained high predictive accuracy over extended follow-up periods, with time-dependent AUCs of 0.950 (95% CI 0.910-0.991), 0.931 (95% CI 0.887-0.976), and 0.909 (95% CI 0.880-0.937) at 3, 5, and 7 years, exceeding multi-clinical variables by 6.8%, 2.9%, and 5.0%, respectively (Fig 3i). Combining PanDerm’s predictions with clinical factors further improved long-term prognostic accuracy in AUCs at 5 and 7 years. PanDerm also consistently outperformed other AI approaches (Fig 3j), showing improvements of 2.3%, 3.0%, and 2.5% at 3, 5, and 7 years respectively. Further details are provided in Methods.
@@
We next evaluated PanDerm’s capability in analyzing whole-body imaging (Total Body Photography, TBP) [2, 3, 47] (Fig 4a). Unlike close-up imaging of individual lesions, TBP enables comprehensive patient-level analysis, particularly for critical melanoma risk factors such as photodamage and nevus count [4, 5, 48]. In a cohort of 480 patients with 196,933 lesions from Australia, PanDerm achieved a weighted F1 score of 0.896 (95% CI 0.879-0.913) for photodamage assessment and an AUROC of 0.983 (95% CI 0.979-0.987) for nevus counting, surpassing all other models (P $<$ 0.05 and P $<$ 0.001, respectively; Fig 4b, c, g). Even with limited training data (10% of the full dataset), PanDerm maintained superior performance (Extended Data Fig 10). In lesion-specific risk stratification, PanDerm also ranked first with an AUROC of 0.705 (95% CI 0.698-0.712) and BACC of 0.657 (95% CI 0.6513-0.663), with all results statistically significant (P $<$ 0.001; Fig 4d, h).
@@
Figure 4: Skin phenotype assessment and melanoma screening using TBP. a. Illustration of PanDerm processing multimodal TBP data (HOP & MYM dataset) for skin examination, phenotype score quantification, risk stratification, and malignant lesion screening. b-c. Class distribution of skin phenotype for solar damage risk (b) and nevus count (c) in datasets. d-e. Class distribution of risk groups and malignant lesions. g. Solar damage risk assessment and nevus counting performance by W_F1 and AUROC. h. Risk prediction performance by AUROC and BACC (Balanced Accuracy). Error bars in g-h show 95% CIs; bar centers represent mean value. Estimates were computed with nonparametric bootstrapping using 1000 bootstrap replicates. P-values calculated with a two-sided t-test. j. Malignant lesion screening performance by sensitivity. Left: using only TBP data. Right: integrating measurement information. The numbers below the bars indicate the recommended suspicious lesion count. k. Number of malignant lesions detected in the test set. f. UMAP plot of PanDerm screening results for test lesions. i. UMAP plot of human screening results for test lesions. l. UMAP plot of PanDerm risk prediction results for test lesions.
@@
In a clinical validation study, PanDerm effectively identified malignant lesions among a large number of benign ones (216 malignant vs. 197,716 benign lesions) from the HOP [49] and MYM [50] cohort (Fig 4e). Using TBP images alone, PanDerm achieved a sensitivity of 0.893, outperforming the next-best model by 4.2% (Fig 4j left). When clinical measurements were available for all models, PanDerm maintained its advantage with a 3.5% higher sensitivity (Fig 4j right), reaching a sensitivity of 0.893. Significantly, it detected malignant lesions in 79 out of 80 patients while reducing unnecessary examinations by 60.8% compared to melanographers (3498 vs. 8913 lesions recommended for detailed examination) (Fig 4j, k; Extended Data Table 17).
@@
We observed that PanDerm’s analysis approach aligned well with established clinical practice, particularly the ”ugly duckling” concept [51] of identifying atypical lesions through comparison with a patient’s other lesions. This was demonstrated through UMAP visualization (Fig 4f), where PanDerm’s feature effectively separated suspicious lesions. The clustering patterns in PanDerm’s risk assessment (Fig 4l) showed correspondence closely with human screening patterns (Fig 4i), illustrating its exceptional performance in malignant lesion screening. Additional details are provided in the Methods section, Extended Data Table 14-17, and Extended Data Fig 11.
@@
We evaluated PanDerm’s performance on skin lesion segmentation using ISIC2018 [52] and HAM10000 [34] datasets. Compared to existing methods including SL-Imagenet, autoSMIM [33], and BATFormer [33], PanDerm achieved significantly higher performance, surpassing the next-best by 3.1% and 1.9% in the Jaccard index on both datasets (P $<$ 0.001; Extended Data Fig 1a, b). PanDerm’s performance was particularly noteworthy in label-limited scenarios, matching the next-best model while using only 5% of the training data (104 and 350 images for ISIC2018 and HAM10000, respectively; Extended Data Fig 1c, d). When compared to MedSAM [53], a medical image segmentation foundation model, PanDerm showed slightly better accuracy (0.5% improvement, P = 0.025 and 0.112; Extended Data Table 20). This is particularly impressive as PanDerm achieves this performance without specialized training for image segmentation. Additionally, PanDerm offers practical advantages in clinical settings, processing images about 4-5 times faster than MedSAM while using less computational resources (Extended Data Table 21). Visual examples and detailed performance metrics are provided in Extended Data Fig 2 and Table 18-21.
Figure 5: Performance of PanDerm in Human-AI Collaborative Assessment of 128 Skin Conditions using Clinical Images. a. Number of readers (n=37) in this study. Each reviewed up to 50 of 200 cases. Readers were divided into two groups based on specialization: The dermatology group (n=20), including dermatology residents and specialists, and the generalist group (n=17), including pre-vocational trainees, general practitioners, nurses, and clinical trial assistants. b. Geographic distribution of readers. c-e. Comparisons without and with PanDerm support across all readers (n=37) for: (c) Top-1 diagnostic assessment score (1-4), (d) Top-3 diagnostic accuracy, and (e) diagnostic confidence score (1-4). P-values in c-e calculated using two-sided paired t-test. f. Change ratio in diagnosis after PanDerm support across specialization groups. g-h. Comparisons without and with PanDerm support stratified by specialization groups for: (g) Top-1 diagnostic assessment score and (h) Top-3 diagnostic accuracy. P-values in g-h calculated using two-sided paired t-test (n=126 classes). i-j. Comparisons without and with PanDerm support stratified by major class types for: (i) Top-1 diagnostic assessment score and (j) Top-3 diagnostic accuracy. P-values in i-j calculated using two-sided paired t-test (n=126 classes).
Reader studies
To assess PanDerm’s clinical applicability, we conducted three reader studies evaluating its capabilities across different aspects and modalities of dermatological diagnosis, as follows.
Early melanoma detection compared with clinicians. To examine PanDerm’s capability in early melanoma detection, we compared it with 12 human reviewers (7 experienced dermatologists and 5 dermatologist trainees) using sequential dermoscopic images from Alfred Hospital [54], featuring multiple follow-up images of the same lesions over time. The study evaluated two key aspects: overall diagnostic accuracy and early melanoma detection capability. In terms of overall accuracy, PanDerm outperformed the average human reviewer by 10.2% and surpassed the best-performing human by 3.6%. For early detection, we assessed the time point of the first suspicious changes detected in sequential images relative to clinical diagnosis and biopsy confirmation. PanDerm demonstrated superior ability in this challenging task, correctly identifying 77.5% (69 out of 89) of melanoma lesions at the first imaging time point, compared to only 32.6% (29 correct diagnoses) for human reviewers (Extended Data Fig 3). Individual dots in the histograms represent the earliest correct diagnosis time points for both PanDerm and human reviewers, visualizing the comparative early detection performance.
@@
Human-AI collaboration for skin cancer diagnosis using dermoscopic images. We evaluated PanDerm’s impact on clinicians’ diagnostic accuracy across seven pigmented lesion classes using dermoscopic images (Extended Data Fig 4a). The study included 41 clinicians with varying levels of competency who evaluated cases both with and without PanDerm’s multi-probability prediction support. PanDerm’s assistance significantly increased overall diagnostic accuracy from 0.69 (95% CI 0.65-0.73) to 0.80 (95% CI 0.76-0.84, P $<$ 0.001; Extended Data Fig 4b). Notably, clinicians with lower competency levels showed the greatest improvement, with accuracy gains of 17% (P = 0.0082) for those with low competency and 12% (P $<$ 0.001) for those with medium competency, while highly competent clinicians demonstrated a 6% improvement (P = 0.039; Extended Data Fig 4c, Extended Data Table 33). Class-specific analysis revealed significant accuracy improvements in 4 of 7 lesion classes (P $<$ 0.05; Extended Data Fig 4d, Extended Data Table 32). For melanoma diagnosis specifically, PanDerm enhanced clinician accuracy from 0.69 (95% CI 0.64-0.74) to 0.83 (95% CI 0.79-0.87, P $<$ 0.001).
Human-AI collaboration for differential diagnosis of 128 skin conditions using clinical photos. We conducted a comprehensive reader study evaluating PanDerm’s diagnostic capabilities across 128 skin conditions using clinical photos. The study included 37 readers from 5 countries (Fig 5b) and comprised two groups(Fig 5a): the dermatology group (n=20; 11 dermatology trainees and 9 specialists) and the generalist group (n=17; 7 pre-vocational trainees, 8 GPs, 1 nurse, and 1 clinical trial assistant). This grouping represents the distinction in specialty training backgrounds between dermatology-trained practitioners and those with general medical training. Each reader assessed up to 50 cases from a 200-case pool, providing their top-3 diagnoses both with and without PanDerm’s assistance. Each reader evaluated up to 50 cases from a 200-case pool, providing their top-3 diagnoses with and without PanDerm’s support. Four experienced dermatologists developed a standardized ontology for condition categorization (Extended Data Fig 5). Performance was assessed primarily using two metrics: a 4-point diagnostic assessment scale for top-1 diagnosis (4: exact ontology match to 1: significant mismatch) and top-3 diagnostic accuracy, with three independent dermatologists scoring and resolving discrepancies through panel review. PanDerm’s assistance significantly improved average top-1 diagnostic scores of all readers from 2.83 to 3.08 (P $<$ 0.001; Fig 5c) and top-3 diagnostic accuracy from 54% to 63.4% (P $<$ 0.001; Fig 5d), while increasing readers’ diagnostic confidence (2.17 to 2.42, P $<$ 0.001; Fig 5e). The impact was particularly pronounced in the generalist group, showing higher diagnosis modification rates (28.6% vs 12.9% in dermatology group; Fig 5f) and greater improvements in both top-1 diagnostic scores (generalist group: +0.45, dermatology group: +0.25; Fig 5g) and top-3 accuracy (generalist group: +16.5%, dermatology group: +10.3%; Fig 5h). Analysis by condition classes demonstrated consistent improvements across inflammatory, neoplastic, and other categories (P $<$ 0.05; Fig 5i, j), with inflammatory conditions showing the largest gains (+0.36 in top-1 diagnostic scores, +14.2% in top-3 accuracy). Further details on the setup, methodology, readers statistics, and datasets of all three reader studies are provided in Methods, Extended Data 6-8 and Extended Data 30-33.
@@ Discussion
Despite significant advances in AI technology, its application in clinical medicine remains fragmented and underutilized. Current AI systems are often restricted to isolated tasks and are unable to address the diverse demands of medical decision-making. This limits AI’s potential in supporting clinicians in disease diagnosis and management. Dermatology, with its complex requirements, including holistic patient assessment, lesion-specific analysis, and potential use of various imaging modalities, serves as an ideal use case for demonstrating AI’s capabilities across multiple interconnected clinical tasks. Success in this domain could pave the way for broader adoption of AI models across healthcare.
@@
To evaluate PanDerm’s clinical utility, we conducted validations across 28 benchmark datasets, spanning comprehensive skin cancer assessment and a diverse set of primary care dermatological conditions. For skin cancer-related assessment, PanDerm outperformed existing models in specialized tasks across various modalities, including risk stratification of lesions, phenotype assessment, detection of lesion changes and malignancy, multi-class cancer diagnosis, lesion segmentation, and metastasis prediction and prognosis. In particular, PanDerm achieved the most results using only 10% of the task-specific training data typically required by existing models, helping alleviate the scarcity of specialist-labeled data in medical AI. In primary care dermatology settings, PanDerm also outperformed comparative models in diagnosing a diverse set of conditions such as inflammatory diseases, infectious conditions, and frequently encountered dermatoses. These capabilities stem from its rich knowledge representation, developed through pretraining on varied dermatological image modalities and conditions, leading to consistent and significant performance improvements across tasks and modalities.
Three reader studies supported these benchmark findings, suggesting PanDerm’s potential to assist clinical practice across different healthcare settings or specialty training backgrounds. In skin cancer diagnosis, PanDerm demonstrated two main capabilities: it helped improve diagnostic accuracy across clinicians of varying competence levels or specialties when used as a support tool, and notably, showed the ability to identify concerning lesions before clinician detection — potentially facilitating earlier intervention when melanomas are thinner and more amenable to treatment. In general dermatology, PanDerm significantly improved readers’ performance in the differential diagnosis of 128 skin conditions, such as inflammatory dermatoses, cutaneous neoplasms, pigmentary disorders, and hair conditions. The improvement appeared more substantial for generalists, such as primary care providers, in evaluating inflammatory conditions — which represent a considerable portion of dermatological consultations in everyday clinical practice. Given the limited access to specialist expertise in primary care settings, where most skin conditions are initially evaluated [56, 57], these findings indicate PanDerm’s potential to help address the need for dermatological expertise across different healthcare settings through both its technical capabilities and clinical applications.
The scaling behavior observed in PanDerm’s performance, particularly in relation to pretraining dataset size, aligns with recent trends in foundation model development [20, 22, 23, 58]. However, achieving this in dermatology required addressing unique challenges in large-scale medical data acquisition and multimodal data integration. Our analysis revealed two key technical insights with broader implications for medical AI development. First, using CLIP [37] as a teacher model for semantic feature alignment achieved superior training data efficiency (Fig 1f), outperforming DINOv2 [39]. This is especially relevant in healthcare, where assembling large-scale datasets remains a significant challenge - for context, DINOv2 typically requires 142 million images, a scale impractical in medical domains. Second, our masked feature reconstruction approach proved particularly effective at capturing subtle diagnostic features in skin images compared to methods like MAE [19], an essential capability for clinical applications. These advantages enabled PanDerm to improve upon both traditional models such as ResNet50 [59] and ViT large [36], and recent generalist medical foundation models like BiomedGPT [42]. While generalist models advance broader biomedical AI development, their pursuit of breadth across medical domains often limits their effectiveness in specialized areas. Our results suggest that specialty-specific foundation models, designed with consideration of clinical workflows and imaging requirements, may offer more practical solutions. This approach could inform AI development across other medical specialties where multiple imaging modalities play crucial diagnostic roles.
Despite promising results, we acknowledge several limitations in our evaluation scope and methodology. While our validation covered approximately 200 skin conditions across major categories (e.g., inflammatory diseases, infections, neoplasms, benign growths, pigmented lesions, and vascular anomalies), this represents only a fraction of known dermatological conditions (over 1,000 diagnoses), and is smaller than some previous studies (e.g., Liu et al [6] with 445 conditions). Specifically, our coverage of rare genetic disorders, complex systemic diseases, and clinical variants remains limited. Regarding model robustness and fairness, while our benchmark evaluations (Extended Data Tables 28-29) demonstrate consistent performance across different settings (anatomical locations, age groups, genders, and skin tones), several constraints exist: the evaluation mainly reflects overall accuracy rather than disease-specific analysis, has varying disease coverage across anatomical locations, and focuses primarily on single imaging modalities. A more comprehensive evaluation framework [60] integrating these aspects will be necessary for further assessing the PanDerml’s robustness. Further, recent studies [61, 62, 63] have highlighted important challenges in dermatological AI systems, particularly in human-AI interactions. Although our evaluations show stable cross-skin tone performance without explicitly balanced training data (as demonstrated necessary by Daneshjou et al [62]), comprehensive bias assessment requires metrics beyond overall accuracy. Groh et al [63] further revealed that equitable standalone performance may not translate to unbiased human-AI collaboration, which is crucial for clinical deployment. To address these limitations, future work should develop standardized protocols for cross-demographic evaluations using more comprehensive fairness metrics and investigate biases in human-AI collaborative settings. International collaborations like ISIC [55] will be crucial for creating representative datasets and establishing robust fairness standards.
In conclusion, PanDerm demonstrates the potential of multimodal specialty-specific foundation models in addressing the diverse clinical needs across specialized and routine clinical practice in dermatology. Through comprehensive pretraining on diverse dermatological images and validation across multiple clinical scenarios, the model showed robust performance across different use cases. Our development approach, combining systematic data curation, advanced self-supervised learning, and rigorous clinical validation, provides a framework for developing medical AI systems that can adapt to varying levels of clinical expertise and healthcare settings. These findings suggest promising directions for developing foundation models in other medical specialties where the integration of diverse imaging modalities and complex clinical workflows is crucial for patient care.